OpenStudy (anonymous):

simplify: 1/x+h - 1/x

OpenStudy (souvik):

is it\[\frac{ 1 }{ x }+h-\frac{ 1 }{ x }\]?

OpenStudy (anonymous):

No its 1 over x+h

OpenStudy (anonymous):

1/x+h-1/x =>x-x-h/(x+h)x =>-h/x^2+hx

OpenStudy (souvik):

\[\frac{ 1 }{ x+h}-\frac{ 1 }{ x }\] \[\frac{ x-(x+h) }{ x(x+h) }\] \[\frac{ -h }{ x(x+h) }\]